OpenStudy (anonymous):

A particle subject to a force vector F(x,y) = (yi-xj) moves clockwise along the arc of a circle with radius 4 centered at the origin, that begins at (-4, 0) and ends at (0, 4) a) Find the work done by vector F.

OpenStudy (dan815):

all them question marks

OpenStudy (dan815):

click edit and remove them

OpenStudy (anonymous):

there? @dan815

OpenStudy (dan815):

k parametrize the curve w=int f.dr

OpenStudy (dan815):

r(t)=<x(t),y(t)> dr/dt=<x'(t),y'(t)> dr=<x'(t),y'(t)> * dt

OpenStudy (dan815):

f=<x,y>=<x(t),y(t) now f.dr is all in one variable and you can integrate over some time bound

OpenStudy (dan815):

in more general case what u are doing here is, u wanna see how much of your force is in the tangent component of your motion, because thats the component that matters wrt to our force

OpenStudy (dan815):

since word=f*d

ganeshie8 (ganeshie8):

you could also use green's theorem by completing the path along x axis

OpenStudy (anonymous):

so (cost, sint) for the parametrization ?

OpenStudy (dan815):

you have to see your bounded area though, because greens theorem will give u integral all over bounded area curve

OpenStudy (dan815):

you could use greens theorem still though by subtracting an area, from the main integral

OpenStudy (anonymous):

so would the integral be from pi to pi/2?

ganeshie8 (ganeshie8):

subtract the work done along x axis, lets work it both ways and compare the answers

OpenStudy (dan815):

curlf = d/dx(-x)-(d/dy)((y)= -2, :( not path independant so u gotta do long way

OpenStudy (dan815):

-2*k^ to be more accurate

OpenStudy (anonymous):

\[-\int\limits_{\pi/2}^{\pi} (sinti-costj) (-sinti+costj)dt\]

OpenStudy (dan815):

ok that works

OpenStudy (dan815):

u can just say pi to pi.2 and get rid of negative sigh no problems

OpenStudy (dan815):

u will see that u have -(sin^2+cos^2) so

OpenStudy (dan815):

you ust have 1 |pi/2 to pi

OpenStudy (dan815):

which is pi/2

OpenStudy (dan815):

wait u set p something wrong

OpenStudy (dan815):

it shudnt be like that becase x=4costheta,y=4sintehta

OpenStudy (dan815):

you will get a factor of 16 that u are missing

OpenStudy (dan815):

so u get pi/2 * 16 = 8pi

OpenStudy (anonymous):

oh yeah i missed the radius 4

OpenStudy (dan815):

happens when u look at megan fox too long

ganeshie8 (ganeshie8):

im getting 16pi by greens thm

OpenStudy (dan815):

or rose bottom

ganeshie8 (ganeshie8):